# On an approach to canonicalizing elliptic Feynman integrals

### v3

@@ 2.2 The construction of the pre-canonical basis with the Legendre normal form
We now propose a set of candidates for the pre-canonical basis, motivated by the three kinds of standard elliptic integrals:
@@
where $i=5,\cdots,n+1$. Note that the number of master integrals is n1n-1. The first and the last candidates involve differentials of the first and second kinds of standard elliptic integrals, while the rest involve the third kind. Plugging these $\phi_{j}^{\text{pre}}$ into Eq. (15), we get the corresponding master integrals $I_{j}^{\text{pre}}$. Note we do not introduce any elliptic integrals into the integrands at this step.
We can now derive the differential equations of the pre-canonical master integrals with respect to $\lambda$ and $e_{i}$ for $i=5,\cdots,n+1$. We find that they are in a linear form:
@@ 3 Examples
Table 1: The examples we will consider are listed in the above table. The first column shows the integral family, with the abbreviated name used in this paper given in brackets. The second column contains the Feynman diagram for the family. Thick, colored lines represent propagators with non-zero invariant masses, and distinct colors represent different masses. The third column gives the number of kinematic scales; the number of independent dimensionless parameters is one less than this value. The fourth column indicates the sectors we consider, where the basis integrals are transformed into canonical form. The number in brackets is the number of master integrals. The number in the bracket is the number of the master integrals. The last column provides references for the canonical bases of the families. Two of these results are new to this work, and their relevant citations are also given.
@@ 3.1 Unequal-mass Sunrise
The first $4$ integrals are in the elliptic top sector, while the remaining $3$ sub-sector integrals are non-elliptic. For simplicity, we take $m_{3}=m_{1}\neq m_{2}$, and then $I_{11200}=I_{21100}$ and $I_{01100}=I_{11000}$ according to symmetric relations. We will study the family in $d=2-2\varepsilon$, where the non-elliptic master integrals are automatically canonical. We can then focus on the top sector.
@@
To write down the canonical basis for the top sector, we just need to plug Eqs. (60) and Eqs. (61) into Eqs. (10), followed by converting the integrands into linear combinations of the master integrals $I_{11200}$, $I_{12100}$ and $I_{11100}$. Accidentally, the sub-sector dependence of these top-sector canonical basis turns out to be $\varepsilon$-factorized automatically. As we will see in later examples, the sub-sector dependence can be non-trivial in more complicated cases, which requires further manipulation.
@@ 3.2 An elliptic sub-sector of $2$-loop $t\bar{t}W$ production
Figure 2: A family contributing to 22-loop $t\bar{t}W$ production Becchetti:2025qlu is shown in the left panel. There are two four-point elliptic sectors, one of which has 4 scales and its canonical basis has been given in Becchetti:2025oyb . The remaining elliptic sector is shown in the right panel, which is more complicated with 5 scales. In the diagrams, thick orange lines represent top quarks with mass $m_{t}$, the curvy lines represent $W$ bosons with mass $m_{W}$, and the thick blue line is an off-shell particle coming from pinching while all the others are massless.
@@ 3.3 An elliptic branch of $2$-loop $W$-pair production via light quark-antiquark annihilation
The maximal cut of the master integral $I_{011100110}$ is given by
@@
Note that this is NOT the same as the derivative in (29), since the latter lives under the maximal cut with no information about sub-sectors. Using this information, we can find “good” combinations of sub-sector integrals that can be added when promoting $\phi_{n-1}^{\rm can}$ to Feynman integrals without cuts. We will employ this trick for the examples in Sec. 3.4 and Sec. 3.5 as well. When this is done (where the derivative is taken with respect to $x$), we find that there are only 3 entries in the differential equations which are not $\varepsilon$-factorized.
@@ 3.4.1 Definition and two-variable construction under the next-to-maximal cut
The first 4 master integrals are in the top sector. They are chosen such that their differential equations have relatively simple dependencies on the (canonical) sub-sector integrals. This will be helpful in transforming these dependencies into the canonical form. The last two master integrals in the first line of Eq. (82) belong to the elliptic triangle family, whose canonical bases are well-studied in the literature Jiang:2023jmk ; Gorges:2023zgv . For the remaining sub-sectors, we construct the canonical basis integrals with ${\rm d}\log$-forms following Chen:2020uyk ; Chen:2022lzr , and the explicit expressions for them are given in App. C. The canonical bases for these non-elliptic sub-sectors were also studied in Ahmed:2024tsg ; Becchetti:2025rrz .
@@
We now just need to plugging the relevant variables into Eq. (10) to obtain $\phi^{\rm can}$. At this point, we would like to point out an interesting phenomenon. The integrands in Eq. (10) involve a complete elliptic integral of the third kind depending on $e_{6}$, for both the top sector and the elliptic triangle sector. However, from the results of Jiang:2023jmk , we have already known that no such integral is needed for the triangle sector. This hints at relations between complete elliptic integrals of the first kind and the third kind. [^8] One may derive similar relations from the sunrise families. In both cases, we find that such relations are associated with the fact that the integrands of the first kind and the third kind can be combined into a ${\rm d}\log$-form. In the current case, we have
@@ 3.5 Non-planar double box B
Our final example is the non-planar double box B family shown in Fig. 5. It shares a lot of similarities with the family in Sec. 3.4, and the canonical basis is studied in Schwanemann:2024kbg . The kinematics is the same as Eqs. (80) and Eqs. (81). The propagators are also the same as Eqs. (79) except that $\chi=0$, so that $D_{1}$ and $D_{3}$ are massless. For this family, we will only consider the top sector and the elliptic triangle sub-sector [^9] Note that there is actually another elliptic curve associated with a different sub-sector, which we do not consider here., which share the same elliptic curve. The starting master integrals are chosen as
@@ Acknowledgements
We would like to thank Ming Lian for collaborations in the early stage of this work. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12375097, 12535003, 12347103, Science Foundation of China University of Petroleum, Beijing (No.2462025YJRC019), and the Fundamental Research Funds for the Central Universities.
Appendix A Construct $\varepsilon$-factorized basis with constant period matrix
@@
Now, the remaining task is to further transform $I_{1}^{\rm pre}$ and $I_{n-1}^{\rm pre}$, to rotate the corresponding $2\times 2$ block of $\bm{A}^{(0)}$ away, which is only $\lambda$ relevant. To do that, we only need to consider the corresponding $2\times 2$ block of the period matrix for $I_{1}^{\rm pre}$ and $I_{n-1}^{\rm pre}$, which is
@@
It is straightforward to combine the transformation above with the pre-canonical basis in Eqs. (48) together with the Möbius transformation in Eq. (2.1), to obtain the corresponding $\varepsilon$-factorized basis at the integrand level. This supports the idea that our workflow can be combined with any algorithms to obtain integrand-level results, and then applied to generic integral families.
@@
Note that $K(\lambda)$ and $F(\sin^{-1}a_{i},\lambda)$ are singular as $c_{1}\to c_{2}$, and they are logarithmically singular in this limit. These singularities do not cause problems for the basis integrals, since they cancel due to the fact that $\phi_{1}^{\rm pre}=\phi_{2}^{\rm pre}$ when $c_{1}\to c_{2}$; see Eq. (49). This cancellation also implies that the first $\varepsilon$-factorized basis integral vanishes in this limit. Similar to the pre-canonical basis, other $\varepsilon$-factorized basis integrals are ${\rm d}\log$-forms asymptotically in the neighborhood of the degenerate points $c_{1}=c_{2}$, as desired. From the demonstration above, we can see the construction in Sec. 2.3 is essentially different from the procedure here from two related perspectives. On the one hand, we introduce $\varepsilon$-dependence in the fiber transformations in the former construction, while the construction here does not. On the other hand, there are only complete elliptic integrals with a unique modulus arguments in the former case, while here we introduce incomplete elliptic integrals and both of complementary moduli arguments are needed. As a consequence, the construction here will introduce more elliptic integrals and undesirable asymptotic properties for the transformation. More relevant discussions can also be found in Frellesvig:2023iwr .
@@ Appendix B Canonical basis with Jacobi normal form
We construct the pre-canonical basis and the corresponding canonical basis with the elliptic curve in Legendre normal form in Sec. 2. From there, we know the explicit dependence on shape of the curve encoded in the parameter $\lambda$ makes our construction simple. There is another special normal form named Jacobi normal form shares similar feature with Legendre normal form. Thus, a natural question is why not try constructing the basis in Jacobi form. In this appendix, we will demonstrate the construction of pre-canonical basis and canonical basis in Jacobi normal form.
@@ B.2.1 The construction of the canonical basis
With $\cal T$$^{\prime}$ from (131) and $\cal T$} from (137), the transformation from pre-canonical basis $\vec{I}^{\rm pre}$ to canonical basis ${\cal T}_{\rm can}$ can be obtained directly. Acting ${\cal T}_{\rm can}$ on Eqs. (125), we obtain a canonical basis in the Jacobi family. The canonical basis and the connection matrix here share similar properties as the ones in Legendre family.
@@ B.2.2 The construction of an $\varepsilon$-factorized basis with constant period matrix
The transformation from Eqs. (154) to the $\varepsilon$-factorized basis is
@@
Similarly, with the transformation, the period matrix is a constant matrix.
@@ B.2.3 From Jacobi family back to the original family
We consider the Möbius transformation from a generic family to the Jacobi family above, which seems to be complicated in general. One may wonder although the transformation is cumbersome for generic cases, are there some special configurations of $c_{i}$’s such that the transformation is simple. Fortunately, the answer is positive. For example, suppose we have the four branch points (the ones related to the shape of the elliptic curve, instead of the ones corresponding to marked points) in a symmetric manner
@@
The Möbius transformation to its Jacobi family is easy to construct since it is just a combination of translation and dialation. To be more explicit, a possible Möbius transformation is
@@
where $i\geq 5$ and $c_{i}$’s correspond to the marked points, and a special case is $a_{\infty}=\infty$. This behavior is as expected since in the transformation $S(z)$ Eq. (149) there is no inversion which is the only transformation that alters the point at the infinity. We can easily read out from Eqs. (150) that the corresponding $k$ is
@@
In order to express the pre-canonical basis in the original family, we also need the relations for the inverse transformation, which can be easily read out from Eqs. (152). The pre-canonical basis in the original family is then straightforward
@@
where the $k$ specifying the Jacobi curve here is the modulus $k^{2}=\lambda$. We can see the twist of the Jacobi family is in a more restricted form, where the marked points $\pm a_{i}$ appear in pairs. Moreover, the number of the basis integrals is twice as many as the original. However, we can always construct the additional pre-canonical basis integrals as ${\rm d}\log$-forms, which disentangles with other ones in the $\varepsilon^{0}$-part of the differential equations. However, the ${\rm d}\log$-forms cannot be transformed back to the original family, while the other ones belong to the original family so they can indeed serve as pre-canonical basis integrals in the original family.
@@ B.3 Application to non-planar double box family
Although the simple pre-canonical basis integrals in Eqs. (154) ask a special original family, it turns out our first non-planar double box family in Sec. 3.4 satisfies the condition. It is important to note that the underlying elliptic curve possesses a unique property: the branch points exhibit an additional symmetry. This symmetry makes the construction in Jacobi form simpler to use.
For the branch points of the single mass case in Sec. 3.4, we can refer back to Eqs. (85). With the Möbius transformation in Eq. (149), the variables $k$ and $a_{5}$ for the elliptic curve in the Jacobi normal form are
@@
Plugging Eqs. (159) into Eqs. (125), one can check this pre-canonical basis enjoys an arguably simpler form than the one obtained in Sec. 3.4, and here algebraic factors only appear in $\phi_{2}^{\rm pre}$. We also note that the pre-canonical basis integrals in Becchetti:2025rrz are very similar to the ones here.

### v4

@@ 2.2 The construction of the pre-canonical basis with the Legendre normal form
We now propose a set of candidates for the pre-canonical basis, motivated by the three kinds of standard elliptic integrals: [^5] Note that given the twist in Eq. (13), the dimension of the twisted cohomology group is n1n-1, which is the number of solutions to the equation $\mathrm{d}u=0$. This corresponds to the number of master integrals in this sector under IBP relations without taking into account possible symmetry relations. This counting is independent of the specific underlying geometry.
@@
where $i=5,\cdots,n+1$. The first and the last candidates involve differentials of the first and second kinds of standard elliptic integrals, while the rest involve the third kind. Plugging these $\phi_{j}^{\text{pre}}$ into Eq. (15), we get the corresponding master integrals $I_{j}^{\text{pre}}$. Note we do not introduce any elliptic integrals into the integrands at this step.
We can now derive the differential equations under maximal cut for the pre-canonical master integrals with respect to $\lambda$ and $e_{i}$ for $i=5,\cdots,n+1$. We find that they are in a linear form:
@@ 3 Examples
Table 1: The examples we will consider are listed in the above table. The first column shows the integral family, with the abbreviated name used in this paper given in brackets. The second column contains the Feynman diagram for the family. Thick, colored lines represent propagators with non-zero invariant masses, and distinct colors represent different masses. The third column gives the number of kinematic scales; the number of independent dimensionless parameters is one less than this value. The fourth column indicates the sectors we consider, where the basis integrals are transformed into canonical form. The number in brackets is the number of master integrals. The last column provides references for the canonical bases of the families. Two of these results are new to this work, and their relevant citations are also given.
@@ 3.1 Unequal-mass Sunrise
The first $4$ integrals are in the elliptic top sector, while the remaining $3$ sub-sector integrals are non-elliptic. We will study the family in $d=2-2\varepsilon$, where the non-elliptic master integrals are automatically canonical. We can then focus on the top sector.
@@
To write down the canonical basis for the top sector, we just need to plug Eqs. (60) and (61) into Eq. (10), followed by converting the integrands into linear combinations of the master integrals $I_{11200}$, $I_{12100}$ and $I_{11100}$. Accidentally, the sub-sector dependence of these top-sector canonical basis turns out to be $\varepsilon$-factorized automatically. As we will see in later examples, the sub-sector dependence can be non-trivial in more complicated cases, which requires further manipulation.
@@ 3.2 An elliptic sub-sector of $2$-loop $t\bar{t}W$ production
Figure 2: A family contributing to 22-loop $t\bar{t}W$ production Becchetti:2025qlu is shown in the left panel. There are two four-point elliptic sectors, one of which has 4 scales and its canonical basis has been given in Becchetti:2025oyb . The remaining elliptic sector is shown in the right panel, which is more complicated with 5 scales. In the diagrams, thick orange lines represent top quarks with mass $m_{t}$, the curvy lines represent $W$ bosons with mass $m_{W}$, and the thick blue line is an off-shell particle coming from pinching the propagator with momentum flow $k_{1}-p_{2}$ while all the others are massless.
@@ 3.3 An elliptic branch of $2$-loop $W$-pair production via light quark-antiquark annihilation
The maximal cut of the master integral $I_{011200110}$ is given by
@@
Note that this is NOT the same as the derivative in (29), since the latter lives under the maximal cut with no information about sub-sectors. In other words, while the canonical forms are uniquely fixed under the maximal cut, their extension to full Feynman integrals is not unique. One has the freedom to add sub-sector integrals which vanish under the maximal cut. These choices directly determine the complexity of the off-diagonal blocks. By using the information from the derivative, we can identify specific “good” combinations of sub-sector integrals to add when promoting $\phi_{n-1}^{\rm can}$ to Feynman integrals without cuts. We will employ this trick for the examples in Sec. 3.4 and Sec. 3.5 as well. When this is done (where the derivative is taken with respect to $x$), we find that there are only 3 entries in the differential equations which are not $\varepsilon$-factorized.
@@ 3.4.1 Definition and two-variable construction under the next-to-maximal cut
The first 4 master integrals are in the top sector. They are chosen such that their differential equations have relatively simple dependencies on the (canonical) sub-sector integrals. This will be helpful in transforming these dependencies into the canonical form. Note that our construction for the top sector integrals under the maximal cut does not depend on the choice, and different choices are the same the under the maximal cut while lead to different sub-sector dependence. The last two master integrals in the first line of Eq. (82) belong to the elliptic triangle family, whose canonical bases are well-studied in the literature Jiang:2023jmk ; Gorges:2023zgv . For the remaining sub-sectors, we construct the canonical basis integrals with ${\rm d}\log$-forms following Chen:2020uyk ; Chen:2022lzr , and the explicit expressions for them are given in App. C. The canonical bases for these non-elliptic sub-sectors were also studied in Ahmed:2024tsg ; Becchetti:2025rrz .
@@
We now just need to plug the relevant variables into Eq. (10) to obtain $\phi^{\rm can}$. At this point, we would like to point out an interesting phenomenon. The integrands in Eq. (10) involve a complete elliptic integral of the third kind depending on $e_{6}$, for both the top sector and the elliptic triangle sector. However, from the results of Jiang:2023jmk , we have already known that no such integral is needed for the triangle sector. This hints at relations between complete elliptic integrals of the first kind and the third kind. [^9] One may derive similar relations from the sunrise families. In both cases, we find that such relations are associated with the fact that the integrands of the first kind and the third kind can be combined into a ${\rm d}\log$-form. In the current case, we have
@@ 3.5 Non-planar double box B
Our final example is the non-planar double box B family shown in Fig. 5. It shares a lot of similarities with the family in Sec. 3.4, and the canonical basis is studied in Schwanemann:2024kbg . The kinematics is the same as Eqs. (80) and (81). The propagators are also the same as Eq. (79) except that $\chi=0$, so that $D_{1}$ and $D_{3}$ are massless. For this family, we will only consider the top sector and the elliptic triangle sub-sector [^10] Note that there is actually another elliptic curve associated with a different sub-sector, which we do not consider here., which share the same elliptic curve. The starting master integrals are chosen as
@@ Acknowledgements
We would like to thank Ming Lian for collaborations in the early stage of this work. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12505094, 12375097, 12535003, 12547104, Science Foundation of China University of Petroleum, Beijing (No.2462025YJRC019), and the Fundamental Research Funds for the Central Universities.
Appendix A Construction of $\varepsilon$-factorized basis with constant period matrix
@@
Now, the remaining task is to further transform $I_{1}^{\rm pre}$ and $I_{n-1}^{\rm pre}$ to remove the corresponding $2\times 2$ block in $\bm{A}^{(0)}$ (23), where the transformation only depends on $\lambda$. To do that, we only need to consider the corresponding $2\times 2$ block of the period matrix for $I_{1}^{\rm pre}$ and $I_{n-1}^{\rm pre}$, which is
@@
It is straightforward to combine the transformation above with the pre-canonical basis in Eq. (48) together with the Möbius transformation in Eq. (2.1), to obtain the corresponding $\varepsilon$-factorized basis at the integrand level. This supports the idea that our workflow can be combined with any algorithms to obtain integrand-level results, and then applied to generic integral families.
@@
Note that $K(\lambda)$ and $F(\sin^{-1}a_{i},\lambda)$ are singular as $c_{1}\to c_{2}$, and they are logarithmically singular in this limit. These singularities do not cause problems for the basis integrals, since they cancel due to the fact that $\phi_{1}^{\rm pre}=\phi_{2}^{\rm pre}$ when $c_{1}\to c_{2}$, see Eq. (49). This cancellation also implies that the first $\varepsilon$-factorized basis integral vanishes in this limit. Similar to the pre-canonical basis, other $\varepsilon$-factorized basis integrals are ${\rm d}\log$-forms asymptotically in the neighborhood of the degenerate points $c_{1}=c_{2}$, as desired.
From the demonstration above, we can see the construction in Sec. 2.3 is essentially different from the procedure here from two related perspectives. On the one hand, we introduce $\varepsilon$-dependence in the fiber transformations in the former construction, while the construction here does not. On the other hand, there are only complete elliptic integrals with a unique modulus arguments in the former case, while here we introduce incomplete elliptic integrals and both of complementary moduli arguments are needed. As a consequence, the construction here will introduce more elliptic integrals and undesirable asymptotic properties for the transformation. More relevant discussions can also be found in Frellesvig:2023iwr .
@@ Appendix B Canonical basis with Jacobi normal form
We construct the pre-canonical basis and the corresponding canonical basis with the elliptic curve in Legendre normal form in Sec. 2. From there, we know the explicit dependence on shape of the curve encoded in the parameter $\lambda$ makes our construction simple. There is another special normal form named Jacobi normal form which shares similar feature with Legendre normal form. Thus, a natural question is why not try constructing the basis in Jacobi form. In this appendix, we will demonstrate the construction of pre-canonical basis and canonical basis in Jacobi normal form.
@@ B.2.1 The construction of the canonical basis
With $\cal T$$^{\prime}$ from Eq. (131) and $\cal T$} from Eq. (137), the transformation from pre-canonical basis $\vec{I}^{\rm pre}$ to canonical basis ${\cal T}_{\rm can}$ can be obtained directly. Acting ${\cal T}_{\rm can}$ on Eq. (125), we obtain a canonical basis in the Jacobi family. The canonical basis and the connection matrix here share similar properties as the ones in Legendre family.
@@ B.2.2 The construction of an $\varepsilon$-factorized basis with constant period matrix
The transformation from Eq. (125) to the $\varepsilon$-factorized basis is
@@
Similarly, with this transformation, the period matrix is a constant matrix.
@@ B.2.3 From Jacobi family back to the original family
We consider the Möbius transformation from a generic family to the Jacobi family above, which seems to be complicated in general. One may wonder although the transformation is cumbersome for generic cases, are there some special configurations of $c_{i}$’s such that the transformation is a simple linear one. Fortunately, the answer is positive. For example, suppose we have the four branch points (the ones related to the shape of the elliptic curve, instead of the ones corresponding to marked points) in a symmetric manner
@@
The Möbius transformation to its Jacobi family is easy to construct since it is just a combination of translation and dilation. To be more explicit, a possible Möbius transformation is
@@
where $i\geq 5$ and $c_{i}$’s correspond to the marked points, and a special case is $a_{\infty}=\infty$. This behavior is as expected since in the transformation $S(z)$ Eq. (149) there is no inversion which is the only transformation that alters the point at the infinity. We can easily read out from Eq. (150) that the corresponding $k$ is
@@
In order to express the pre-canonical basis in the original family, we also need the relations for the inverse transformation, which can be easily read out from Eq. (152). The pre-canonical basis in the original family is then straightforward
@@
where the $k$ specifying the Jacobi curve here is the modulus $k^{2}=\lambda$. We can see the twist of the Jacobi family is different from Eq. (122), in that the marked points $\pm a_{i}$ now appear in pairs. As a result, the number of the basis integrals is twice as many as the original family. To deal with that, besides pre-canonical basis integrals in Eq. (125), we can construct additional ones as ${\rm d}\log$-forms. These additional integrals disentangle with those in Eq. (125) in the $\varepsilon^{0}$-part of the differential equations. These additional ${\rm d}\log$-forms do not belong to the original family, and are intrinsic in the Jacobi family (158). To see that, it is easy to show that when inverse-transforming from $t$ back to the variable $z$, these ${\rm d}\log$-forms necessarily involve square-roots of the form (157), which do not appear in the twist of the original family. In contrast, the integrals in Eq. (125) would not have such square-roots under the inverse transformation, so they can indeed serve as pre-canonical basis integrals in the original family.
@@ B.3 Application to non-planar double box family
Although the simple pre-canonical basis integrals in Eq. (154) ask a special original family, it turns out our first non-planar double box family in Sec. 3.4 satisfies the condition. It is important to note that the underlying elliptic curve possesses a unique property: the branch points exhibit an additional symmetry. This symmetry makes the construction in Jacobi form simpler to use.
For the branch points of the single mass case in Sec. 3.4, we can refer back to Eq. (85). With the Möbius transformation in Eq. (149), the variables $k$ and $a_{5}$ for the elliptic curve in the Jacobi normal form are
@@
Plugging Eq. (159) into Eq. (125), one can check this pre-canonical basis enjoys an arguably simpler form than the one obtained in Sec. 3.4, and here algebraic factors only appear in $\phi_{2}^{\rm pre}$. We also note that the pre-canonical basis integrals in Becchetti:2025rrz are very similar to the ones here.